# GALD-SE: Guided Anisotropic Lightweight Diffusion for Efficient Speech Enhancement

### v1

@@ Abstract
Speech enhancement is designed to enhance the intelligibility and quality of speech across diverse noise conditions. Recently, diffusion model has gained lots of attention in speech enhancement area, achieving competitive results. Current diffusion-based methods blur the initial signal with isotropic Gaussian noise and recover clean speech from the prior. However, these methods often suffer from a substantial computational burden. We argue that the inefficiency stems from the oversight that speech enhancement is not purely a generative task; it primarily involves noise reduction and completion of missing information, while the clean clues in the original mixture do not need to be regenerated. In this paper, we propose a method that introduces noise with anisotropic guidance during the diffusion process, allowing the neural network to focus on clean clues within noisy recordings. This approach is robust against various types of noise interference and speech distortion, and significantly reduces the computational load. Experiments demonstrate that the proposed method achieves state-of-the-art results with only approximately 4.5 million parameters, significantly fewer than those required by other diffusion methods. This effectively narrows the gap in model size between diffusion-based and predictive speech enhancement approaches. Additionally, the proposed method performs well in very noisy scenarios, demonstrating its potential for applications in highly challenging environments.
@@ I Introduction
Diffusion-based methods operate through a two-phase process: gradually adding noise to an initial state in the forward process, and then iteratively sampling along the reverse process to generate targeted samples. In the field of speech enhancement, the spectrogram of noisy speech frequently serves as a component of the prior for the diffusion process and as an input to the neural network for reverse sampling [9], [10], [11], [13], or exclusively the latter [17, 18, 22]. The corresponding diffusion process can be interpreted in a discrete [29] or continuous [9] way, while the latter permits the use of numerical stochastic differential equation (SDE) solvers to generate estimated speech [25]. Richter et al. [9] proposed an SDE-based method that introduces a novel drift term. In this way, the mean state of the process is transformed from clean speech into its noisy counterpart during the forward process and reversed during sampling. Two main approaches have been explored to improve upon this scheme: (i) Carefully designing new interpolation techniques, such as the Brownian bridge [15, 16] and variance-preserving interpolation methods [13, 14], which have demonstrated significant performance improvements; and (ii) Integrating diffusion models with predictive methods [12, 20, 23]. The latter approach, incorporating a predictive model, can sometimes reduce the computational complexity of the original diffusion method. However, even the smallest version requires approximately 27 million parameters to achieve satisfactory results [23], making diffusion-based methods highly computationally intensive. We argue that this burden stems from the inefficiency of the diffusion process. Current diffusion scheme doesn’t effectively make use of the clean clues in the mixture and isotropically blur them during the forward process. This blurring distorts the original clean clues, forcing the neural network to re-estimate them during the reverse process, thereby increasing the computational load.
Based on these insights, we propose GALD-SE, a Guided Anisotropic Lightweight Diffusion-based Speech Enhancement method, which incorporates anisotropic guidance in the diffusion process. The guidance effectively preserves the integrity of clean clues within noisy mixtures, enabling the process to directly access these clues at any timestep. The formulation of anisotropic diffusion process is built upon the foundational work presented in [27, 28]. Experimental results demonstrate that our diffusion-based method achieves state-of-the-art performance while utilizing only 4.5 million parameters, which is less than 20% of the parameter count in the smallest existing diffusion model [23]. Ablation analysis and comparison confirm the efficiency and superior quality of generated speech across various performance metrics.
@@ II Proposed Method
The system overview of our method is depicted in Fig. 2. An anisotropic noise guidance is derived from a coarse magnitude estimation network. This guidance directs the diffusion process to blur each time-frequency (T-F) bin with specific intensity. During each step of the reverse process, the diffusion network takes the current state, anisotropic guidance, and the noisy speech as inputs. It then computes an estimate of the initial state, which corresponds to the clean spectrogram. This estimated initial state is then used to derive the state at the previous timestep, generating the final result through multiple iterations. By implementing anisotropic guidance to minimize unnecessary diffusion, we enhance both the efficiency and performance. The rest of this section is organized as follows: First, we explain the modifications made to the original diffusion process. Next, we provide the calculation of anisotropic guidance. Finally, we describe the training and sampling strategies of our method.
@@ II-A Stochastic Process
We denote the complex STFT feature vectors of clean and noisy speech as $\mathbf{x}_{0}$ and $\mathbf{y}$, respectively. These vectors reside in $\mathbb{C}^{d}$, where $d=KF$, representing the coefficients of a flattened complex spectrogram. In the typical “ResShift” formulation [27], during the forward process, the mean of the state gradually transitions from $\mathbf{x}_{0}$ to $\mathbf{y}$ by adding isotropic noise at each step. The conditional distribution of the state at time $t$, given the state at time $t-1$, the initial state $\mathbf{x}_{0}$, and the noisy feature $\mathbf{y}$, is given by:
@@
where $\boldsymbol{\sigma}\in\mathbb{R}^{KF\times KF}$ is the anisotropic noise guidance matrix, designed to preserve the clean clues within noisy mixture. This matrix is diagonal, resulting in Gaussian noise that is independent across different T-F bins. During the forward process, the real and imaginary components of the complex features within each T-F bin undergo independent diffusion with the same variance, which is decided by the estimated proportion of noise in that bin. An illustrative example of the anisotropic process is depicted in Fig. 2(a). Suppose that in the original mixture, the 2nd, 3rd, and 5th frequency bins contain a smaller proportion of noise and should therefore be blurred less, while the 1st, 4th, and 6th bins contain a larger proportion of noise and should undergo more blurring. Consequently, for the 2nd, 3rd and 5th bins, the Gaussian noise added at each step is relatively small, keeping the values at these spectral positions nearly unchanged. In contrast, for the other bins, the Gaussian noise added at each timestep is much larger. In this way, the diffusion process can selectively blur T-F bins with varying intensities during the forward process, thereby preventing the introduction of excessive noise to bins that predominantly contain clean speech components.
Although the anisotropic guidance introduces varying intensities of noise to different bins, the overall mean of the process remains unaffected due to the zero-mean nature of the Gaussian noise. Consequently, following the derivation in [27], we can directly sample $\mathbf{x}_{t}$ from the initial state $\mathbf{x}_{0}$ as follows:
@@
where $\beta_{t}=\frac{\alpha_{t}}{\overline{\alpha}_{t}}$ and $\overline{\boldsymbol{\sigma}}_{t}^{2}=\kappa^{2}\beta_{t}(1-\beta_{t}){\boldsymbol{\sigma}}^{2}$. The calculation of the anisotropic guidance $\boldsymbol{\sigma}$ will be detailed in the following section.
@@ II-B Anisotropic Guidance
We developed anisotropic guidance specifically to blur the noise structure while minimally altering the clean components. To achieve this, we consider the ideal ratio mask (IRM), which represents the proportion of magnitude of clean components present in each T-F bin. By utilizing a lightweight UNet model, denoted as $\mathbf{g}_{\delta}(\mathbf{y})$, which is adapted from [30] and parameterized by $\delta$, we input the noisy feature $\mathbf{y}$ to obtain a clean magnitude estimation. Subsequently, we derive a coarse IRM, denoted as $\tilde{\mathbf{M}}_{0}$, residing in $\mathbb{R}^{K\times F}$:
$\tilde{\mathbf{M}}_{0}=\mathbf{g}_{\delta}(\mathbf{y})/\left\|\mathbf{y}\right\|_{2}$ (5)
We then define the guidance matrix $\boldsymbol{\sigma}$ to be diagonal, with its elements obtained as follows:
$\boldsymbol{\sigma}=\operatorname{diag}(\operatorname{vec}(\mathbf{1}-\tilde{\mathbf{M}}_{0}))$ (6)
where $\mathbf{1}$ is a matrix in $\mathbb{R}^{K\times F}$ with all elements set to 1. It is evident that the guidance $\boldsymbol{\sigma}$ is derived from the proportion of noise present in each T-F bin. The higher the noise ratio in a bin, the greater the corresponding value of $\sigma$ is in that bin. In this way, we encourage that the noise in the origin mixture loses its original structure during the forward process, while the clean structure remains nearly intact. This guidance simplifies the model’s task of distinguishing clean speech from the mixture and reduces the risk of receiving misleading clues from the noise parts.
Figure 1: System overview.
@@
We further give an visualization for the impact of anisotropic noise guidance as in Fig. 2(b). The top row displays the noisy spectrogram, followed by the prior sampling states without and with guidance, respectively. The second row indicates the corresponding clean speech and the enhanced results. We observe that with anisotropic guidance, the T-F structure of the noise is significantly blurred, while the initial state more effectively preserves clean speech clues compared to its counterpart without guidance. This preservation facilitates the model’s estimation of the original clean spectrogram. Consequently, the enhancement result with anisotropic guidance retains more speech details than those without guidance.
@@ II-C Training and Sampling
At training stage, we sample a state at an arbitrary time $t$ as described in (3), incorporating guidance from (6). The diffusion network, denoted as $\mathbf{f}_{\theta}(\mathbf{x}_{t},\mathbf{y},\boldsymbol{\sigma},t)$, is parameterized by $\theta$ and processes inputs including the current state $\mathbf{x}_{t}$, noisy feature $\mathbf{y}$, anisotropic guidance $\boldsymbol{\sigma}$, and the current timestep $t$. We design the model to estimate the initial state $\mathbf{x}_{0}$, as the presence of anisotropic guidance complicates the direct estimation of Gaussian noise. This model, along with the coarse magnitude estimation network $\mathbf{g}_{\delta}(\mathbf{y})$ discussed in Section II-B, are simultaneously trained using a mean square error criterion, leading to the following optimization objective:
$\min_{\boldsymbol{\theta},\boldsymbol{\delta}}\sum_{t}\left[\left\|\mathbf{f}_{\theta}\left(\mathbf{x}_{t},\mathbf{y},\boldsymbol{\sigma},t\right)-\mathbf{x}_{0}\right\|_{2}^{2}+\left\|\mathbf{g}_{\delta}(\mathbf{y})-\left\|\mathbf{x}_{0}\right\|_{2}\right\|_{2}^{2}\right]$ (7)
At sampling stage, we first employ (5), (6) to derive the anisotropic guidance $\boldsymbol{\sigma}$, and then get $\overline{\boldsymbol{\sigma}}_{T}$ based on (4). To initiate the reverse sampling procedure, we sample the initial state as follows:
@@
At each reverse step $t$, we derive the previous state $t-1$ as follows:
@@
where $\beta_{t}$ and $\overline{\boldsymbol{\sigma}}_{t}$ retain their definitions as in (4). By applying this reverse sampling process, we iteratively derive the final estimation $\mathbf{\tilde{x}}_{0}$ from the initial state $\mathbf{x}_{T}$.
@@ III Experiments
III-A Datasets and Evaluation Metrics
We use DNS-Challenge [31] and VoiceBank+Demand (VBD) [32] as the verified datasets. For the DNS corpus, we generated 1,000,000 pairs for training and 300 pairs for validation, each lasting 6 seconds, with signal-to-noise ratios (SNR) ranging from -15 to 25 dB. To construct the test set, we randomly selected 1000 English read speech clips, distinct from those in the training set, and 1000 noise clips, each with a duration of 6 seconds. These clips were mixed at various SNR levels ranging from -15 dB to 15 dB, divided into three intervals. The VBD corpus, widely used as a benchmark dataset in speech enhancement, consists of a training set with 28 speakers subjected to 8 noise types across 5 SNR levels, and a test set featuring 2 speakers with unseen noise types.
We choose the following widely used metrics to evaluate the performance: 1) PESQ: the Perceptional Evaluation of Speech Quality [33]; 2) ESTOI: the Extended Short-Time Objective Intelligibility [34]; 3) SI-SNR: Scale-Invariant Signal-to-Noise Ratio; 4) CSIG, CBAK, COVL: the speech quality (CSIG), background noise quality (CBAK) and the overall quality (COVL) of the audio [35] which reflects the mean opinion scores; 5) DNSMOS: A non-intrusive perceptual objective speech quality metric to evaluate noise suppressors [36].
TABLE I: Performance comparison of different methods on DNS Dataset
@@ III-A Datasets and Evaluation Metrics
GALD-SE (Ours) D+G 4.5M 10 1.57 2.20 2.96 2.25 0.62 0.81 0.89 0.77 6.44 13.07 18.71 12.74 3.73 3.92 3.94 3.86
@@ III-B Hyperparameters and Training Settings
For the stochastic process, we set the number of diffusion steps $T=10$ and the noise variance strength $k=0.19$ in (2). The noise schedule $\overline{\alpha}_{t}$ is bounded between $\overline{\alpha}_{1}=0.001$ and $\overline{\alpha}_{T}=0.999$. A geometric schedule is adopted for intermediate steps: $\sqrt{\overline{\alpha}_{t}}=\sqrt{\overline{\alpha}_{1}}\times\left(\sqrt{\frac{\overline{\alpha}_{T}}{\overline{\alpha}_{1}}}\right)^{\frac{t-1}{T-1}}$. We adapted the diffusion network architecture from SGMSE+ [9]. The feature dimension is set to 32, resulting in 3.6 million parameters for the diffusion model. The coarse magnitude estimation network, adapted from [30], uses only a magnitude mask decoder, contributing 0.9 million parameters. Consequently, the entire system comprises 4.5 million parameters.
For the STFT configuration, we employ a 510-point FFT with a hop size of 128 samples. Following [9], we compress the resulting spectrum using an exponential factor of 0.5 and then scale it by a factor of 0.15. For training, we set the batch size to 15 and use a learning rate of $10^{-4}$. The model is trained for 500 epochs on the VBD corpus and 30 epochs on the DNS corpus.

### v5

@@ Abstract
Speech enhancement is designed to enhance the intelligibility and quality of speech across diverse noise conditions. Recently, diffusion model has gained lots of attention in speech enhancement area, achieving competitive results. Current diffusion-based methods blur the signal with isotropic Gaussian noise and recover clean speech from the prior. However, these methods often suffer from a substantial computational burden. We argue that the computational inefficiency partially stems from the oversight that speech enhancement is not purely a generative task; it primarily involves noise reduction and completion of missing information, while the clean clues in the original mixture do not need to be regenerated. In this paper, we propose a method that introduces noise with anisotropic guidance during the diffusion process, allowing the neural network to preserve clean clues within noisy recordings. This approach substantially reduces computational complexity while exhibiting robustness against various forms of noise and speech distortion. Experiments demonstrate that the proposed method achieves state-of-the-art results with only approximately 4.5 million parameters, a number significantly lower than that required by other diffusion methods. This effectively narrows the model size disparity between diffusion-based and predictive speech enhancement approaches. Additionally, the proposed method performs well in very noisy scenarios, demonstrating its potential for applications in highly challenging environments. The code is available at https://github.com/wangchengzhong/GALDSE.
@@ I Introduction
Diffusion-based methods operate through a two-phase process: gradually adding noise to an initial state in the forward process, and then iteratively sampling along the reverse process to generate targeted samples. In the field of speech enhancement, the spectrogram of noisy speech frequently serves as a component of the prior for the diffusion process and as an input to the neural network for reverse sampling [9], [10], [11], [13], or exclusively the latter [18, 22]. The corresponding diffusion process can be interpreted in a discrete [29] or continuous [9] way, while the latter permits the use of numerical stochastic differential equation (SDE) solvers to generate target speech [26]. Richter et al. [9] proposed an SDE-based method that introduces a novel drift term. In this way, the mean state of the process is transformed from clean speech into its noisy counterpart during the forward process and reversed during sampling. Two main approaches have been explored to improve upon this scheme: (i) Carefully designing new interpolation techniques, such as the Brownian bridge [15, 16], variance-preserving interpolation methods [13, 14] and Schrödinger bridge [24], which have demonstrated significant performance improvements; and (ii) Integrating diffusion models with predictive methods [12, 20, 23]. These methods can sometimes reduce the computational complexity [23, 24], however, they’re still highly computational-intensive. We argue that this burden stems from the inefficiency of the diffusion process when applied to speech enhancement tasks. The uniform corruption fails to preserve the clean speech features in the mixture, resulting in a large distributional gap between diffusion states and clean speech, which may increase computational costs during inference.
Based on these insights, we propose GALD-SE, a Guided Anisotropic Lightweight Diffusion-based Speech Enhancement method, which incorporates anisotropic noise guidance in the diffusion process. The guidance effectively preserves the integrity of clean clues within noisy mixtures, enabling the process to directly access these clues at any timestep. The formulation of anisotropic diffusion is built upon the foundational work presented in [28]. Experimental results demonstrate that our method achieves performance comparable to state-of-the-art while utilizing only 4.5 million parameters. Ablation analysis and comparison confirm the efficiency and superior quality of generated speech across various performance metrics.
@@ II Proposed Method
We begin with an overview of our proposed system, as illustrated in Fig. 2. Our system enhances speech by leveraging the anisotropic diffusion process, with the anisotropic guidance derived from a coarse mask estimation. The system comprises three main components: (i) a coarse mask estimation network for deriving guidance, (ii) a guided anisotropic diffusion process for selectively blurring noisy components, and (iii) a diffusion network for iterative refinement. The enhancement procedure begins by estimating the anisotropic guidance and sampling a prior state. It then proceeds through our guided reverse diffusion process facilitated by the diffusion network, ultimately yielding the enhanced clean speech. While the core network architectures are adapted from existing models [9, 30], our key innovation lies in the anisotropic design of the diffusion process, which significantly reduces the computational load while achieving superior performance. The following subsections detail our main contributions: modifications to the original diffusion process, computation of anisotropic guidance, and specialized training and sampling strategies.
@@ II-A Stochastic Process
We denote the complex STFT feature vectors of clean and noisy speech as $\mathbf{x}_{0}$ and $\mathbf{y}$ respectively, both residing in $\mathbb{C}^{d}$ where $d=KF$, representing the coefficients of a flattened complex spectrogram. In the typical “ResShift” formulation [28], during the forward process, the mean of the state gradually transitions from $\mathbf{x}_{0}$ to $\mathbf{y}$ by adding isotropic noise at each step. The conditional distribution of the state at time $t$, given the state at time $t-1$, the initial state $\mathbf{x}_{0}$, and the noisy feature $\mathbf{y}$, is given by:
@@
where $\boldsymbol{\sigma}\in\mathbb{R}^{KF\times KF}$ is the anisotropic noise guidance matrix, designed to preserve the clean clues within noisy mixture. This matrix is diagonal, resulting in Gaussian noise that is independent across different T-F bins. During the forward process, the real and imaginary components of the complex features within each T-F bin undergo independent diffusion with the same variance, which is decided by the estimated proportion of noise in that bin. Fig. 2(a) illustrates the anisotropic process. For clarity, we focus on the real component of the state; the imaginary component follows a similar pattern. Suppose that in the noisy feature vector, indices 2, 3, and 5 contain a smaller proportion of noise and should therefore be blurred less, while indices 1, 4, and 6 contain a larger proportion of noise and should undergo more blurring. Consequently, for indices 2, 3, and 5, the Gaussian noise added at each step is relatively small, keeping the values at these positions nearly unchanged. In contrast, for the other indices, the Gaussian noise added at each timestep is much larger. In this way, the diffusion process can selectively blur T-F bins with varying intensities during the forward process, thereby preventing the introduction of excessive noise to bins that predominantly contain clean speech components.
Although the anisotropic guidance introduces varying intensities of noise to different bins, the overall mean of the process remains unaffected due to the zero-mean nature of the Gaussian noise. Consequently, following [28], we can directly sample $\mathbf{x}_{t}$ from the initial state $\mathbf{x}_{0}$ as follows:
@@
where $\beta_{t}=\frac{\alpha_{t}}{\overline{\alpha}_{t}}$ and $\overline{\boldsymbol{\sigma}}_{t}^{2}=\kappa^{2}\beta_{t}(1-\beta_{t}){\boldsymbol{\sigma}}^{2}$. The calculation of the anisotropic guidance $\boldsymbol{\sigma}$ will be detailed in the following subsection.
@@ II-B Anisotropic Guidance
We developed anisotropic guidance specifically to blur the noise structure while minimally altering the clean components. CMEN, denoted as $\mathbf{g}_{\delta}(\mathbf{y})$, is applied to derive this guidance. It is a lightweight UNet model adapted from [30], which employs a representative structure commonly used in speech enhancement. The network takes $\mathbf{y}$ as input to estimate the truncated phase-sensitive mask ${\mathbf{M}}\in\mathbb{R}^{K\times F}$ [31], which is bounded within $[0,1]$:
$M_{k,f}=\text{clip}_{[0,1]}\left(\cos\theta_{k,f}\cdot\frac{|x_{0}|_{k,f}}{|y|_{k,f}}\right)$ (5)
where $\theta$ denotes the phase difference between $x_{0}$ and $y$. The phase-sensitive design aligns with our complex-domain diffusion process, providing a coarse estimation of speech proportion in each T-F bin. We use the estimated mask $\tilde{\mathbf{M}}:=\mathbf{g}_{\delta}(\mathbf{y})$ to define the guidance matrix $\boldsymbol{\sigma}$, which is diagonal, with its elements obtained as follows:
$\boldsymbol{\sigma}=\operatorname{diag}(\operatorname{vec}(\mathbf{1}-\tilde{\mathbf{M}}))$ (6)
where $\mathbf{1}$ is a matrix in $\mathbb{R}^{K\times F}$ with all elements set to 1. The higher the noise ratio in a bin, the greater the corresponding value of $\sigma$ is in that bin. In this way, we encourage that the noise in the origin mixture loses its original structure during the forward process, while the clean structure remains nearly intact.
Figure 1: System Overview. An anisotropic guidance is derived from a coarse mask estimation network and subsequently applied to the diffusion process. The prior state is sampled using noisy speech and guidance, after which the clean output is obtained by iteratively executing the reverse process, facilitated by the diffusion network. For detailed model structures, refer to [9, 30].
@@
We further give an visualization for the impact of anisotropic noise guidance as in Fig. 2(b). The top row displays the noisy spectrogram, followed by the prior sampling states without and with guidance, respectively. The second row indicates the corresponding clean speech and the enhanced results. We observe that with anisotropic guidance, the T-F structure of the noise is significantly blurred, while the clean speech clues are preserved more effectively compared to its counterpart without guidance. This preservation facilitates the model’s estimation of the original clean spectrogram. Consequently, the enhancement result with anisotropic guidance retains more speech details than those without guidance.
@@ II-C Training and Sampling
We follow the ResShift [28] formulation to perform training and sampling, albeit with the incorporation of anisotropic guidance. The diffusion network, denoted as $\mathbf{f}_{\theta}(\mathbf{x}_{t},\mathbf{y},\boldsymbol{\sigma},t)$, is parameterized by $\theta$ and processes inputs including the current state $\mathbf{x}_{t}$, noisy feature $\mathbf{y}$, anisotropic guidance $\boldsymbol{\sigma}$, and the current timestep $t$. The guidance $\boldsymbol{\sigma}$ is derived from the CMEN through (6), allowing the diffusion network to adapt to potential estimation errors in CMEN. At training stage, we design the diffusion model to estimate the initial state $\mathbf{x}_{0}$, as the presence of anisotropic guidance complicates the direct estimation of Gaussian noise. The CMEN is trained using the phase-sensitive spectrum approximation described in [31]. These two components are simultaneously trained using a mean square error criterion, leading to the following optimization objective:
$\min_{\theta,\delta}\sum_{t}\left[\|\mathbf{f}_{\theta}(\mathbf{x}_{t},\mathbf{y},\boldsymbol{\sigma},t)-\mathbf{x}_{0}\|_{2}^{2}+\|\mathbf{g}_{\delta}(\mathbf{y})\odot\mathbf{y}-\mathbf{x}_{0}\|_{2}^{2}\right]$ (7)
where $\odot$ denotes element-wise multiplication. It should be noticed that gradients from the diffusion network are not backpropagated to the CMEN, allowing each component to focus on its specific task.
At sampling stage, we first employ (6) to derive the anisotropic guidance $\boldsymbol{\sigma}$, and then get $\overline{\boldsymbol{\sigma}}_{T}$ based on (4). To initiate the reverse sampling procedure, we sample the prior state as follows:
@@
At each reverse step $t$, we derive the previous state $\mathbf{x}_{t-1}$ using the estimated initial state $\tilde{\mathbf{x}}_{0}=\mathbf{f}_{\theta}(\mathbf{x}_{t},\mathbf{y},\boldsymbol{\sigma},t)$:
@@
where $\beta_{t}$ and $\overline{\boldsymbol{\sigma}}_{t}$ retain their definitions as in (4). By applying this reverse sampling process, we iteratively derive the final estimation $\mathbf{\tilde{x}}_{0,\text{final}}$ from the prior state $\mathbf{x}_{T}$.
@@ III Experiments
TABLE I: Performance comparison of different methods on DNS Simu Dataset
@@
GALD-SE (Ours) D+G 4.5M 6 1.61 2.23 3.02 2.29 0.63 0.83 0.89 0.78 6.83 13.42 19.32 13.19 3.69 3.83 3.92 3.81
III-A Datasets and Evaluation Metrics
We use DNS-Challenge [32] and VoiceBank+Demand (VBD) [33] as the verified datasets. For the DNS corpus, we generated 1,000,000 pairs for training and 300 pairs for validation, each lasting 6 seconds, with signal-to-noise ratios (SNR) ranging from -15 to 25 dB. To construct the simulated test set, we randomly selected 1000 English read speech clips, distinct from those in the training set, and 1000 noise clips, each with a duration of 6 seconds. These clips were mixed at various SNR levels ranging from -15 dB to 15 dB, divided into three intervals. The VBD corpus, widely used as a benchmark dataset in speech enhancement, consists of a training set with 28 speakers subjected to 8 noise types across 5 SNR levels, and a test set featuring 2 speakers with unseen noise types.
We choose the following widely used metrics to evaluate the performance: 1) PESQ: the Perceptional Evaluation of Speech Quality [34]; 2) ESTOI: the Extended Short-Time Objective Intelligibility [35]; 3) SI-SNR: Scale-Invariant Signal-to-Noise Ratio; 4) CSIG, CBAK, COVL: the speech quality (CSIG), background noise quality (CBAK) and the overall quality (COVL) of audio [36]; 5) DNSMOS: A non-intrusive perceptual objective speech quality metric to evaluate noise suppressors [37].
@@ III-B Hyperparameters and Training Settings
For the stochastic process, we set the number of diffusion steps $T=6$ and the noise variance strength $\kappa=0.5$ in (2). The noise schedule $\overline{\alpha}_{t}$ is bounded between $\overline{\alpha}_{1}=0.001$ and $\overline{\alpha}_{T}=0.999$. A non-uniform geometric schedule is adopted for intermediate steps: ${\overline{\alpha}_{t}}={\overline{\alpha}_{1}}\times\left({{\overline{\alpha}_{T}}/{\overline{\alpha}_{1}}}\right)^{\left(\frac{t-1}{T-1}\right)^{p}}$, where $p=0.3$. We adapted the diffusion network architecture from [9]. As in Fig. 1, the first-step feature dimension is set to 32, resulting in 3.6 million parameters for the diffusion model. The CMEN, adapted from [30], uses only a mask decoder, contributing 0.9 million parameters. Consequently, the entire system comprises 4.5 million parameters.
For the STFT configuration, we employ a 510-point FFT with a hop size of 128 samples. We compress the resulting spectrum using an exponential factor of 0.5 and then scale it by a factor of 0.5. For training, we set the batch size to 15 and use a learning rate of $10^{-4}$. The model is trained for 1000 epochs on the VBD corpus and 60 epochs on the DNS corpus.